OpenStudy (anonymous):

Solve. x^2 – x = 20.

OpenStudy (mathmate):

By completing squares: \[x^2 – x = 20\] \[(x–\frac{1}{2})^2 = 20 + \frac{1}{4}\] \[x-\frac{1}{2} = \pm \sqrt{(\frac{81}{4})} = \pm \frac{9}{2}\] \[x=\frac{1}{2} \pm \frac{9}{2} = 5\ or -4\]